OpenStudy (anonymous):

find the inverse of f(x)=2x-8

OpenStudy (anonymous):

x=2y-8 2y=x+8 y=(x+8)/2

OpenStudy (anonymous):

function says a) multiply by 2 b) subtract 8 inverse says a) add 8 b) divide by 2 so it is \[f^{-1}(x)=\frac{x+8}{2}\] like if you were solving \[2x-8=20\]

OpenStudy (anonymous):

thanks alot